OpenStudy (kitkat16):

AB passes through A(-3,0) and B(-6,5) what is the equationof the line that passes through the origin and parallel to AB

OpenStudy (kitkat16):

so i found the slope between the two points m=5/-3 m=-5/3 so would the equation be 5x-3y=0

OpenStudy (kitkat16):

@ganeshie8 can you check this for me?

ganeshie8 (ganeshie8):

slope is correct,

ganeshie8 (ganeshie8):

equation thru origin is : y = mx y = (-5/3)x 5x + 3y = 0

OpenStudy (kitkat16):

thank you!!!